### v1

@@ Abstract
We examine the location characteristics of a conditional selective confidence interval based on the polyhedral method. This interval is constructed from the distribution of a test statistic conditional upon the event of statistical significance. In the case of a one-sided test, the behavior of the interval varies depending on whether the parameter is highly significant or only marginally significant. When the parameter is highly significant, the interval is similar to the usual confidence interval derived without considering selection. However, when the parameter is only marginally significant, the interval falls into an extreme range and deviates greatly from the estimated value of the parameter. In contrast, an interval conditional on two-sided significance does not yield extreme results, although it may exclude the estimated parameter value.
@@
Methods for selective inference provide valid statistical inference on parameters when the estimated model or the parameters of interest are selected based on the observed data. Conventional inference assumes that the model or parameters are fixed in advance. Empirical practice often diverges from this ideal, potentially invalidating conventional methods. For example, standard confidence intervals for parameters that are chosen based on the observed data may not have the desired coverage. Data-dependence may be introduced by formal methods for model selection such as the Lasso (Lee et al., 2016) or by more informal approaches of data-mining. An example of informal data-mining is a $p$-hacking scenario where a researcher estimates many parameters and then crafts a narrative highlighting the few that are statistically significant.
@@
Our setting is the “file drawer problem” (Rosenthal, 1979), that is, the situation where only statistically significant estimates are presented. The main result is on inference on a parameter $\theta$ that is found significant by a one-sided $t$-test. Let $X$ denote the corresponding $t$-statistic and $\bar{x}$ the critical value. Suppose that we observe $X=x^{\text{obs}}$. The parameter is significant if $x^{\text{obs}}\geq\bar{x}$. The polyhedral method for selective inference on $\theta$ is based on the distribution of $X$ conditional on observing $X\geq\bar{x}$, which has the distribution function $F(a;\theta,\bar{x}):=\Pr(X\leq a\mid X\geq\bar{x})$. Note that $F$ depends on $\theta$ because of the dependence of the distribution of $X$ on it. The conditional confidence interval is found by evaluating the distribution function at the observed $t$-statistic $x^{\text{obs}}$ and then inverting $\theta\mapsto F(x^{obs};\theta,\bar{x})$.
@@
Our main result is for situations where $x^{\text{obs}}$ only barely exceeds $\bar{x}$, a case often referred to as “marginal significance”. For this case, we show that the conditional confidence interval is very wide and located far away from the estimated value of the parameter. Specifically, we show that even the upper bound of the interval diverges to $\infty$- ∞ as $x^{\text{obs}}\downarrow\bar{x}$. Therefore, at the margin of significance, the confidence interval contains only extremely negative numbers, a range of numbers that may be a-priori implausible.
@@
Our results are closely related to Kivaranovic & Leeb (2021), who show that the expected length of a confidence interval constructed by the polyhedral method is infinite when the conditioning set is bounded either from below or from above. Indeed, our main result immediately follows from their Lemma A.3. However, our focus differs from theirs. They consider the expected length of the confidence interval. We complement their results by identifying precisely when (that is, in what kind of samples) the confidence interval exhibits extreme behavior. Whereas Kivaranovic & Leeb (2021) focus on the length of the confidence interval, our focus is on its location. We show that the conditional confidence interval for a marginally significant parameter is located far away from its estimate. This result applies more broadly than our file drawer problem, since it builds on Lemma A.3 of Kivaranovic & Leeb (2021) which holds generally for truncated normal distributions. In particular, when the conditioning set is bounded from either below or above and the observed parameter is at the margin of the conditioning set, the conditional selective confidence intervals are located in extreme ranges.
As far as we know, this “location problem” of selection-robust inference on marginally significant parameters has not been previously discussed in the literature. It is of great practical importance, given that empirical $t$-statistics often cluster close to common critical values, as documented by Gerber & Malhotra (2008), Brodeur et al. (2016), and Brodeur et al. (2020). This clustering has raised concerns about potential $p$-hacking and led to marginally significant results being viewed with skepticism. Selection-robust inference methods may be able to address some of the concerns about marginally significant results. However, the “location problem” implies that the difference between robust estimates and the original (conventional) estimates may be extreme.
A parameter estimate can be considered “highly significant” when its $t$-statistic far exceeds the critical value. In this case, the conditional selective confidence interval is almost identical to the conventional interval. This feature has already been documented in the literature. For instance, Andrews et al. (2024) discuss statistical inference for determining the best welfare-improving policy and show that the confidence interval based on the polyhedral method is nearly identical to the conventional one when the best policy is unambiguously determined.
For moderate cases, where the parameter estimate is neither “marginally” nor “highly” significant, the conditional selective confidence interval provides non-trivial insights. It is wider than the conventional interval, but still informative enough to provide a useful range of parameter values. The increased length intuitively reflects additional statistical uncertainty when the data it is used both for selection and inference.
We show that the problem of extreme confidence intervals at the “margin of significance” does not appear when statistical significance is determined by a two-sided test, that is, a test that rejects if $\lvert x^{\text{obs}}\rvert>\bar{x}$. In this case, the limiting interval as $x^{\text{obs}}\downarrow\bar{x}$ is located around zero and has a relatively short length. This is already suggested by the result of Kivaranovic & Leeb (2021) that the expected length of the interval is finite in this case. A perhaps surprising result of our analysis is that, even though the interval is “well-behaved” in the sense that it does not diverge, it may still exclude the conventional estimate of the parameter.
@@
Our arguments are entirely mechanical and do not rely on the underlying distribution of the data. Regardless of the data’s true distribution or the parameter’s actual value, if the observed value of the $t$-statistic is close to the critical value, the problem identified in this paper will arise. This aspect of our results starkly contrasts other studies that highlight potential issues with selective inference methods. For instance, the result in Kivaranovic & Leeb (2021) on the expected length relies on distributional assumptions to evaluate the expectation. Our result does not need any distributional assumptions.
@@
The remainder of the paper is organized as follows. Section 1 introduces the settings. The problems are illustrated via a numerical example in Section 2. Section 3 presents the main analytical results. We discuss the two-sided significance cases in Section 4. Section 5 concludes.
1 Settings
@@
We consider the file drawer problem from Rosenthal (1979) of reporting only statistically significant effects. The selective conditional inference procedure based on the polyhedral approach Lee et al. (2016) is introduced.
@@
We are interested in a parameter $\theta$ for which an estimator $\hat{\theta}$ is available. Assume that the estimator $\hat{\theta}$ is unbiased and Gaussian: $\hat{\theta}\sim N(\theta,\sigma^{2})$. For simplicity, we assume that the variance $\sigma^{2}$ is known. In this case, it is without loss of generality to normalize $\sigma^{2}=1$ and $\hat{\theta}$ is identical to its $t$-statistic $X=X/\sigma$. Let $x^{\text{obs}}$ denote the observed value of $X$. First, we discuss the one-sided significance, where an effect is significant if $x^{\text{obs}}>\bar{x}$ for a pre-specified critical value $\bar{x}$. The two-sided significance case is discussed in Section 4.
@@
We follow the polyhedral approach given in Lee et al. (2016). It is based on the conditional distribution of $X$ given that it exceeds the critical value. Under the Gaussianity assumption, the conditional distribution function is given by
$F(a;\theta,\bar{x}):=$ $\Pr(X\leq a\mid X\geq\bar{x})$
$=$ $\frac{\Phi(a-\theta)-\Phi(\bar{x}-\theta)}{1-\Phi(\bar{x}-\theta)}.$
@@
$p=F(x^{\text{obs}};\theta,\bar{x}),$ (1)
for given $p\in(0,1)$, $x^{\text{obs}}$, and $\bar{x}$. The solution exists by Lemma A.2 in Kivaranovic & Leeb (2021) and is strictly decreasing in $p$. The dependence of $\theta(p)$ on $x^{\text{obs}}$ is crucial for the results in this paper, even though it is kept implicit in the notation.
Various statistics for the selective conditional inference correspond to $\theta(p)$ with specific choices of $p$. The conditionally median unbiased estimator is given by
@@
A $100(1-\alpha)$conditional confidence interval is defined as
@@
For example, $[\theta(0.975),\theta(0.025)]$ is the 95% confidence interval.
To verify that the confidence interval has the desired conditional coverage probability, note that
$\Pr(\theta\in[\theta(1-\alpha/2),\theta(\alpha/2)]\mid X\geq\bar{x})$
$=$ $\Pr(\alpha/2\leq F(X,\theta,\bar{x})\leq 1-\alpha/2\mid X\geq\bar{x})=1-\alpha,$
where the last equality holds because the Rosenblatt (1952) transformation implies that, conditional on $X\geq\bar{x}$, $F(X,\theta,\bar{x})$ is uniformly distributed on the unit interval.
2 A numerical illustration
We now present a numerical illustration of the issues discussed in this note for the case of one-sided significance testing. The critical value is set to $\bar{x}=1.64$, corresponding a nominal level of 5%. We set the targeted (conditional) coverage of the confidence intervals to 95%.
@@ 2 A numerical illustration
NOTE: Conditional inference after one-sided testing with critical value $\bar{x}=1.64$. The figure plots the median-unbiased estimator (solid curve) and the bounds $\theta(0.975)$ and $\theta(0.025)$ of the 95% conditional confidence interval (bold dashed curves). The thin dashed line is the 45-degree line, representing the unconditional estimate. The thin dash-dotted lines give the usual confidence interval.
@@
Figure 1 plots the median-unbiased estimator and the conditional confidence interval as a function of the observed $t$-statistic $x^{\text{obs}}$. The upper dashed curve represents the upper bound $\theta(0.025)$ of the confidence interval and the lower dashed curve represents the lower bound $\theta(0.975)$. The solid curve represents the median unbiased estimator $\hat{\theta}_{MU}$. The dashed line is the 45-degree line. Since we have normalized $X=\hat{\theta}$, it gives the unconditional estimate $\hat{\theta}$. The two parallel dash-dotted lines represent the conventional confidence interval $x^{\text{obs}}\pm 1.96$.
@@
Figure 1 reveals several noteworthy insights. When the estimated parameter is highly significant, i.e., when $x^{\text{obs}}$ is large, selective inference yields almost identical results to conventional inference that ignores the selection problem. In this case, the median-unbiased estimator is close to the conventional parameter estimate ($\hat{\theta}_{MU}\approx\hat{\theta}$) and the conditional confidence interval approximates the traditional equal-tailed interval ($\theta(0.975)\approx x^{\text{obs}}-1.96$ and $\theta(0.025)\approx x^{\text{obs}}+1.96$).
@@
When the observed $t$-statistic is of moderate size, conditional inference results differ from the conventional inference. The median-unbiased estimator is lower than the conventional estimate. The conditional confidence interval is wider than the conventional one. Its lower bound is below the lower bound of the standard interval, whereas the upper bounds of the two intervals are fairly similar. The downward adjustments reflect correction for selection bias. Significant effects are positively selected since overestimated effects are more likely to clear the threshold of significance. The wider interval reflects an increase in statistical uncertainty when the data is used for both parameter selection and inference.
@@
In the case of marginal significance, that is, when $x^{\text{obs}}$ is close to the threshold $\bar{x}$, the downward corrections are quite extreme. As $x^{\text{obs}}$ approaches $\bar{x}$, the upper bound of the conditional confidence interval drops below the critical value $\bar{x}$ and eventually below zero. This property of conditional inference may seem counterintuitive. Significantly positive parameter estimates are often interpreted as providing a “statistical guarantee” of a positive population parameter. This perspective seems incompatible with inferring a negative range of parameter values. However, the guarantees given by significance testing are unconditional and do not contradict the inference results based on the conditional distribution. Still, from a practical perspective, a confidence interval that gives a range of negative values may be deemed implausible and one can simply state that the inference results are uninformative.
@@
The numerical results in Figure 1 are easily analytically verified. For example, the median-unbiased estimator is smaller than the critical value $\bar{x}$ if $\theta(0.5)<\bar{x}$ and the upper-bound of the confidence interval is smaller than $\bar{x}$ if $\theta(0.025)<\bar{x}$. The solution $\theta(p)$ depends on the observed $t$-statistic $x^{\text{obs}}$. To find the range of $x^{\text{obs}}$ such that $\theta(p)<\bar{x}$, solve the inequality
$p>F(x^{\text{obs}};\bar{x},\bar{x})=\frac{\Phi(x^{\text{obs}}-\bar{x})-0.5}{0.5},$
@@
$x^{\text{obs}}<\Phi^{-1}(0.5p+0.5)+\bar{x}.$
Thus, unless $x^{\text{obs}}$ exceeds $\bar{x}$ by more than $\Phi^{-1}(0.5p+0.5)$, $\theta(p)$ is smaller than $\bar{x}$. For example, if $x^{obs}<\Phi^{-1}(0.5\times 0.025+0.5)+\bar{x}\approx 0.031+\bar{x}$, the upper bound of the 95% confidence interval is less than $\bar{x}$.
@@
$x^{\text{obs}}<\Phi^{-1}((1-\Phi(\bar{x}))p+\Phi(\bar{x})).$
Thus, unless $x^{\text{obs}}$ exceeds $\Phi^{-1}((1-\Phi(\bar{x}))p+\Phi(\bar{x}))$, $\theta(p)$ is negative. In particular, the upper bound of a 95% confidence interval is negative when $x^{\text{obs}}<\Phi^{-1}(0.05\times 0.025+0.95)\approx 1.66$. This means that the estimate $\hat{\theta}$ is significantly positive at the selection stage, but significantly negative at the inference stage. This apparent contradiction can be reconciled by noting that the selection stage is based on the unconditional distribution of $\hat{\theta}$, whereas the inference stage is based on its conditional distribution.

### v2

@@ Abstract
We examine the location properties of a conditional selective confidence interval constructed via the polyhedral method. The interval is derived from the distribution of a test statistic conditional on the event of statistical significance. For a one-sided test, its behavior depends on whether the parameter is highly or only marginally significant. In the highly significant case, the interval closely resembles the conventional confidence interval that ignores selection. By contrast, when the parameter is only marginally significant, the interval may shift far to the left of zero, potentially excluding all a priori plausible parameter values. This “location problem” does not arise if significance is determined by a two-sided test or by a one-sided test with randomized response (e.g., data carving).
@@
Methods for selective inference provide valid statistical inference on parameters when the estimated model or the parameters of interest are selected based on the observed data. Conventional inference methods assume a fixed model structure and require pre-specification of the parameters of interest before analysis. Empirical practice often departs from this ideal, potentially invalidating conventional methods. For example, standard confidence intervals for parameters that are chosen based on the observed data may not have the desired coverage. Data-dependence may be introduced by formal methods for model selection such as the Lasso (Lee et al., 2016), or from more informal practices of data-mining. An example of informal data-mining is a $p$-hacking scenario where a researcher estimates many parameters and then crafts a narrative highlighting the few that are statistically significant.
@@
Our setting is the “file drawer problem” (Rosenthal, 1979), that is, the situation where only statistically significant estimates are presented. The main result is on inference on a parameter $\theta$ that is found significant by a one-sided $t$-test. Let $X$ denote the corresponding $t$-statistic and $c$ the critical value. Suppose that we observe $X=x^{\text{obs}}$. The parameter is significant if $x^{\text{obs}}\geq c$. The polyhedral method for selective inference on $\theta$ is based on the distribution of $X$ conditional on observing $X\geq c$, which has the distribution function $F(a;\theta,c):=\Pr(X\leq a\mid X\geq c)$. Note that $F$ depends on $\theta$ because of the dependence of the distribution of $X$ on it. The conditional confidence interval is found by evaluating the distribution function at the observed $t$-statistic $x^{\text{obs}}$ and then inverting $\theta\mapsto F(x^{obs};\theta,c)$.
@@
Our main result is for situations where $x^{\text{obs}}$ only barely exceeds $c$, a case often referred to as “marginal significance”. For this case, we show that the conditional confidence interval is very wide and located far away from any reasonable parameter values. Specifically, we show that even the upper bound of the interval diverges to -$\infty$ as $x^{\text{obs}}\downarrow c$. Therefore, at the margin of significance, the confidence interval contains only highly negative values that may be a-priori implausible.
@@
Our results are closely related to Kivaranovic & Leeb (2021), who show that the expected length of a confidence interval constructed by the polyhedral method is infinite when the conditioning set is bounded either from below or from above. Indeed, our main result immediately follows from their Lemma A.3. However, our focus differs from theirs. They consider the expected length of the confidence interval. We complement their results by identifying precisely when (that is, in what kind of samples) the confidence interval exhibits extreme behavior. Whereas Kivaranovic & Leeb (2021) focus on the length of the confidence interval, our focus is on its location. We show that the conditional confidence interval for a marginally significant parameter can be located in an extreme range of the real line, far to the left of zero. In particular, a researcher who expects the parameter to lie within a certain “plausible range” may find that all plausible values to be excluded from the conditional confidence interval.
This “location problem” applies more broadly than our file drawer problem, since it builds on Lemma A.3 of Kivaranovic & Leeb (2021) which holds generally for truncated normal distributions. In particular, when the conditioning set is bounded from either below or above and the observed parameter is at the margin of the conditioning set, the conditional selective confidence intervals are located in extreme ranges.
@@
Our arguments are entirely mechanical and do not rely on the underlying distribution of the data. Regardless of the data’s true distribution or the parameter’s true value, if the observed value of the $t$-statistic is close to the critical value, the problem identified in this paper will arise. This aspect of our results starkly contrasts other studies that highlight potential issues with selective inference methods. For instance, the result in Kivaranovic & Leeb (2021) on the expected length relies on distributional assumptions to evaluate the expectation. Our result does not need any distributional assumptions.
As far as we know, the location problem has not been previously discussed in the literature. The prevalence of “marginally significant” results in empirical work (Gerber & Malhotra, 2008; Brodeur et al., 2016, 2020) suggests that it has great practical importance. Using conditional inference to account for the selectivity of significant results leaves researchers vulnerable to the location problem, potentially preventing them from inferring plausible parameter values. As we discuss below, researchers can avoid this issue by a careful choice of selection mechanism.
Our analysis of polyhedral inference considers also the cases of a moderately or highly significant estimate in addition to the case of marginal significance.
An estimate is “highly significant” if its $t$-statistic far exceeds the critical value. In such cases, the conditional selective confidence interval is nearly indistinguishable from the conventional interval. This property has already been noted in the literature. For example, Andrews et al. (2024) study inference for identifying the most welfare-improving policy and show that, when the best policy is unambiguously determined, the confidence interval based on the polyhedral method is almost identical to the conventional one.
For moderate cases, where the parameter estimate is neither “marginally” nor “highly” significant, the conditional selective confidence interval reveals non-trivial information about the population parameter. It is wider than the conventional interval, but still informative enough to provide a useful range of parameter values. The increased length intuitively reflects additional statistical uncertainty when the data is used both for selection and inference.
We discuss two modifications to the selection rule in the file drawer problem that eliminate the location problem. The suitability of these modifications depends on the empirical context.
Firstly, we consider selection with a randomized response (Fithian et al., 2017; Tian & Taylor, 2018; Panigrahi et al., 2021; Panigrahi & Taylor, 2023). Conditional inference under this selection rule has been shown to yield confidence intervals with bounded expected lengths (Kivaranovic & Leeb, 2024). We find that such methods avoid the location problem.
Secondly, we replace the one-sided $t$-test by a two-sided $t$-test. With this modification, polyhedral inference does not exhibit the location problem. This is already suggested by the result of Kivaranovic & Leeb (2021) that the expected length of the interval is finite in this case.
@@
The remainder of the paper is organized as follows. Section 1 introduces the settings and illustrates the problems with a numerical example. Section 2 presents the main analytical results. We then discuss alternative scenarios in which the problems can be avoided. Specifically, inferences with randomized response are addressed in Section 3.1, and the two-sided significance cases are explored in Section 3.2. Finally, Section 4 concludes the discussion.
1 Setting
@@
Our setting is the file drawer problem from Rosenthal (1979) of reporting only statistically significant effects. We consider conditional inference based on the polyhedral approach form Lee et al. (2016).
@@
We are interested in a parameter $\theta$ for which an estimator $\hat{\theta}$ is available. Assume that the estimator $\hat{\theta}$ is unbiased and Gaussian: $\hat{\theta}\sim N(\theta,\sigma^{2})$. For simplicity, we assume that the variance $\sigma^{2}$ is known. In this case, it is without loss of generality to normalize $\sigma^{2}=1$ and $\hat{\theta}$ is identical to its $t$-statistic $X=X/\sigma$. Let $x^{\text{obs}}$ denote the observed value of $X$. First, we discuss the one-sided significance, where an effect is significant if $x^{\text{obs}}\geq c$ for a pre-specified critical value $c$. The two-sided significance case is discussed in Section 3.2.
@@
The polyhedral approach is based on the conditional distribution of the $t$-statistic $X$ given that it exceeds the critical value. Under Gaussianity the corresponding conditional distribution function is given by
$F(a;\theta,c):=$ $\Pr(X\leq a\mid X\geq c)$ (1)
$=$ $\frac{\Phi(a-\theta)-\Phi(c-\theta)}{1-\Phi(c-\theta)}.$
@@
$p=F(x^{\text{obs}};\theta,c)$
for given $p\in(0,1)$, $x^{\text{obs}}$, and $c$. The solution exists by Lemma A.2 in Kivaranovic & Leeb (2021) and is strictly decreasing in $p$. The dependence of $\theta(p)$ on $x^{\text{obs}}$ is crucial for the results in this paper, even though it is kept implicit in the notation.
Based on $\theta(p)$ we now define a conditionally media-unbiased estimator and a conditional confidence set. The conditionally median-unbiased estimator is given by
@@
A $100(1-\alpha)$% conditional confidence interval is defined as
@@
For example, $\theta(0.975),\theta(0.025)]$ is the 95% confidence interval.
The validity of the confidence interval is verified by the following:
$\Pr(\theta\in[\theta(1-\alpha/2),\theta(\alpha/2)]\mid X\geq c)$
$=$ $\Pr(\alpha/2\leq F(X,\theta,c)\leq 1-\alpha/2\mid X\geq c)=1-\alpha,$
which relies on the fact that, conditional on $X\geq c$, $F(X,\theta,c)$ is uniformly distributed on the unit interval.
To provide numerical evidence of the key issue discussed in this note, we set $c=1.64$, corresponding to a nominal level of 5% for one-sided significance testing. We set the targeted conditional coverage of the confidence interval to 95%.
@@
NOTE: Conditional inference after one-sided testing with critical value $c=1.64$. The figure plots the median-unbiased estimator (solid curve) and the bounds $\theta(0.975)$ and $\theta(0.025)$ of the 95% conditional confidence interval (dashed curves). The thin dashed line is the 45-degree line, representing the unconditional estimate. The shaded region represents the conventional unconditional 95% confidence interval.
@@
Figure 1 plots the median-unbiased estimator and the conditional confidence interval as a function of the observed $t$-statistic $x^{\text{obs}}$. The upper dashed curve represents the upper bound $\theta(0.025)$ of the confidence interval and the lower dashed curve represents the lower bound $\theta(0.975)$. The solid curve represents the median unbiased estimator $\hat{\theta}_{MU}$. The thin dashed line is the 45-degree line. Since we have normalized $X=\hat{\theta}$, it gives the unconditional estimate $\hat{\theta}$. The shaded region represents the conventional confidence interval $x^{\text{obs}}\pm 1.96$.
@@
Figure 1 reveals several noteworthy insights. When the estimated parameter is highly significant, that is, when $x^{\text{obs}}$ is large, selective inference and conventional inference yield almost identical results. In this case, the median-unbiased estimator is close to the conventional parameter estimate ($\hat{\theta}_{MU}\approx\hat{\theta}$) and the conditional confidence interval approximates the traditional equal-tailed interval, that is, $\theta(0.975)\approx x^{\text{obs}}-1.96$ and $\theta(0.025)\approx x^{\text{obs}}+1.96$.
@@
When the observed $t$-statistic is of moderate size, conditional inference results differ from the conventional inference. The median-unbiased estimator lies below the conventional estimate, and the conditional confidence interval is wider than its conventional counterpart. Its lower bound falls below the lower bound of the standard interval, whereas the upper bounds of the two intervals remain fairly similar. These downward adjustments reflect a correction for selection bias: significant effects are positively selected because overestimated effects are more likely to exceed the threshold of significance. The wider interval reflects increased statistical uncertainty when the same data are used for both parameter selection and inference.
@@
For “marginally significant” results, the downward correction can be quite extreme. As $x^{\text{obs}}$ approaches $c$ from above, the upper bound of the conditional confidence interval falls below the critical value $c$ and eventually below zero. Inferring negative values for a “significantly positive” parameter is not the contradiction it may appear: the significance test controls the unconditional error, whereas the conditional confidence interval gives a conditional guarantee. From a practical perspective, however, a confidence interval that includes large negative values may be deemed implausible, and an empirical researcher may reasonably conclude that the inference results are uninformative.
@@
The numerical results in Figure 1 are easily analytically verified. The solution $\theta(p)$ depends on the observed $t$-statistic $x^{\text{obs}}$. To find the range of $x^{\text{obs}}$ such that $\theta(p)<c$, solve the inequality
$p>F(x^{\text{obs}};c,c)=\frac{\Phi(x^{\text{obs}}-c)-0.5}{0.5},$
@@
$x^{\text{obs}}<\Phi^{-1}(0.5p+0.5)+c.$
Thus, unless $x^{\text{obs}}$ exceeds $c$ by more than $\Phi^{-1}(0.5p+0.5)$, $\theta(p)$ is smaller than $c$. For example, if $x^{obs}<\Phi^{-1}(0.5\times 0.025+0.5)+c\approx 0.031+c$, the upper bound of the 95% confidence interval is less than $c$.
@@
$x^{\text{obs}}<\Phi^{-1}((1-\Phi(c))p+\Phi(c)).$
Thus, unless $x^{\text{obs}}$ exceeds $\Phi^{-1}((1-\Phi(c))p+\Phi(c))$, $\theta(p)$ is negative. In particular, the upper bound of a 95% confidence interval is negative when $x^{\text{obs}}<\Phi^{-1}(0.05\times 0.025+0.95)\approx 1.66$. This means that the estimate $\hat{\theta}$ is significantly positive at the selection stage, but significantly negative at the inference stage. This apparent contradiction can be reconciled by noting that the selection stage is based on the unconditional distribution of $\hat{\theta}$, whereas the inference stage is based on its conditional distribution.